OpenStudy (anonymous):

What vaule of B makes this equation true 3/8 * 8/B equals 1 A. -3 B. 0 C. 3 D. 8 I will give you a medal and be your fan!!!

OpenStudy (anonymous):

I think it's C

OpenStudy (anonymous):

You can solve this either with the understanding that multiplying a fraction by it's reciprocal will give the value of 1 (this 3/8 * 8/3 = 1) or by solving for the variable: 3/8 * 8/b = 1 3 * 8 / 8 * b = 1 24 / 8b = 1 3/b = 1 3 * b/b = 1 * b 3 = b

OpenStudy (anonymous):

You are correct, by the way.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!! Xo!